OpenStudy (phoenixfire):

Double Integrals Find the volume of the solid \(z=(x^2 + y^2)-1\) between the two plates z=0 and z=1 I'm not sure how to start. I've graphed the solid, but that hasn't helped me figure out what limits to use.

OpenStudy (e.mccormick):

The limits would be the z axis change.

OpenStudy (e.mccormick):

Then you rotate around y, with a floating radius of x.

OpenStudy (e.mccormick):

That would be one way to do it.

OpenStudy (anonymous):

Is that z = x^2 + y^2 - 1? @PhoenixFire

OpenStudy (phoenixfire):

Yeah, I don't know why they put the ( ) there. it's the same as \[z=x^2+y^2-1\]

OpenStudy (e.mccormick):

It would be a shift on the z axis. You can still reduce it to a single axis problem, which was what I was thinking of.

OpenStudy (phoenixfire):

@e.mccormick I'm not quite sure how you would do that.

OpenStudy (e.mccormick):

Or I think you can... hehe. I have been known to be wrong!

OpenStudy (phoenixfire):

@Hoa Yes! Give us some direction if you know how to do it.

OpenStudy (e.mccormick):

@Hoa Please do, I am checking my statement! And I may be completely off here.

OpenStudy (zugzwang):

lol hoa dont bring real analysis into integrals confusing!

OpenStudy (phoenixfire):

@Hoa I think polar coordinates might be the way to do this.

OpenStudy (zugzwang):

no, i'm still studying this

OpenStudy (zugzwang):

im just a chess player lol :D but okay let's see...

OpenStudy (zugzwang):

polar coords okay?

OpenStudy (zugzwang):

it looks less menacing that way \[\large z = r^2 - 1\]

OpenStudy (phoenixfire):

Maybe it is a triple integral. I thought it would be double.

OpenStudy (zugzwang):

guys u still there?

OpenStudy (phoenixfire):

Yup, just looking through my textbook to see if maybe it is a Triple. But keep going with your solution

OpenStudy (zugzwang):

i dont have one. :(

OpenStudy (zugzwang):

maybe it does have to be in triple integrals...

OpenStudy (phoenixfire):

No I actually do not have answers. It's part of my assignment. We haven't learned triple integrals yet, so maybe I'm getting ahead of myself.

OpenStudy (phoenixfire):

Alright guys, thanks for trying. I'm going to close this question and do some research into it. I'll come back and ask again in a few days if it's not sorted.

OpenStudy (zugzwang):

i have an idea

OpenStudy (zugzwang):

how about \[\Large \int\limits_{\theta = 0}^{2\pi}\int\limits_{r=1}^{\sqrt 2}(r^2-1) \ rdrd\theta\]

OpenStudy (zugzwang):

if r goes from 1 to square root of 2, then z would go from 0 to 1.

OpenStudy (zugzwang):

im sorry i didn't see besides you probably deleted it

OpenStudy (zugzwang):

yes we meet lol hi im Kevin Nguyen and you are...?

OpenStudy (zugzwang):

american, vietnamese dad

OpenStudy (zugzwang):

no i dont speak tieng viet

OpenStudy (zugzwang):

you gave one, i gave it again only wait for his reaction are you a high school student?

OpenStudy (zugzwang):

big brother/sister lol

OpenStudy (zugzwang):

dont know what that means

OpenStudy (zugzwang):

maybe I will you better not have cussed at me tho lol

OpenStudy (zugzwang):

No. I'm a college freshman.

OpenStudy (phoenixfire):

@zugzwang That will give me the region  Right? Which isn't what we want.  This has to be a triple integral. I'll see if I understand it next week when we do triple integral.